OpenStudy (anonymous):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable z is eliminated in the first and third equations, then the second and third equations. 5x + 3y – 2z = 16 8x + y – 2z = 8 3x + y + 2z = 0

OpenStudy (turingtest):

-3x+2y=8 8x+y-2z=8 11x+2y=8

OpenStudy (anonymous):

thank you (: i just worked it out on my ownn !

OpenStudy (turingtest):

even better! welcome